You are in charge of a banquet for your family reunion. The cost is $5.50 per children plate and $12.00 per adult plate. The total number of people is 546. The total bill for the food is $5115.50. Using a for adult plate and c for child plate, write an equation to represent the number of adults and children present.
Answer by checkley79(3341)

5.50c+12.00a=$5,115.50
a+c=546 or a=546-c
5.50c+12.00(546-c)=5,115.50
5.50c+6,552-12c=5,115.50
-6.5c=5,115.50-6,552
-6.5c=-,436.50
c=-1,436.50/-6.5
c=221 ans. for the number of children.
a=546-21=325 ans. for the number of adults.
Proof:
5.50*221+12.00*325=5,115.50
1,215.50+3,900=5,115.50
5,115.50=5,155.50